A uniform bar weighing 60 N has a length of 4 m and is supported at its end by strings. A load W=200N is placed at some point on the bar such that it is in equilibrium when the tension in the string at the left end is 80N. (a) Locate the position of W, (b) find the tension in the string at the right end.


1
Expert's answer
2021-07-12T12:14:36-0400
"T_2=200+60-80=180\ N\\60(2)+200(x)=4(180)\\x=3\ m"